# Finding 3D Positions of Distant Objects from Noisy Camera Movement and Semantic Segmentation Sequences

### v1

@@ IV Experiments and Results
IV-A Simulation
Simulation offers a possibility to study the method extensively by enabling testing in an unlimited number of different scenarios varying by, for example, the camera trajectories, noise levels, and observation distances. Here, the simulation results are presented in an order of increasing complexity. Since the number of possible variations is infinite, we tried to limit the results to scenarios which could offer the most insight into the method’s performance in the expected real-world tasks as well as its robustness to the different noise sources. The simulation results are presented in Table I.
The camera intrinsic parameters were set as $f_{x}=1200$, $f_{y}=1200$, $c_{x}=960$, and $c_{y}=540$, to imitate the camera used in the empirical test. The discretisation caused by the camera pixels was taken into account in all simulations.
The simulated perpendicular observation trajectory presents the optimal and simplest target observation scenario. Without noise, this presents an exceedingly optimal situation, and we mainly used it to confirm that the filter performs as expected and to analyse the minimum uncertainty and errors which could theoretically be obtained when locating visual targets using the particle filter. We performed these tests with a single 100 by 100 by 100 metre target, with the target position in metres, $tp=(x,y,z)^{\top}=(500,-200,2000)^{\top}$. The trajectory of the camera positions in metres, $cp$, started from $cp=(0,0,0)^{\top}$ and continued linearly to $cp=(1000,0,0)^{\top}$. The camera angle remained stationary in the simulation, and it was defined such that when the target and camera positions’ x-coordinates were equal, the target was projected in the horizontal middle axis of the camera frame. The pitch angle of the camera was set such that the horizon was in the middle of the camera frame’s vertical axis.
Even with perfectly noiseless segmentation and an optimal trajectory, the camera requires some translation for the position estimate to converge to the correct position. However, it’s worth noting that the range of RMSEs is still only between approximately 15% and 2% relative error in the target position estimate for a target at a two-kilometre distance.
The main simulation results are collected in Table I with the corresponding simulation experiments explained in the following paragraphs. The convergence of the filter in the corresponding experiments with single and multiple targets are visualised in Figure LABEL:fig:singletarget and Figure LABEL:fig:multitarget, respectively.

### v2

@@ IV Experiments and Results
Our experiments demonstrate that the proposed particle filter can estimate distant target positions as effectively as multi-view triangulation. The evaluation was done using a simulation, visualised in Figure LABEL:fig:singletarget, and two drone-captured image segmentation sequences of a telecommunication mast shown in Figure LABEL:fig:rw_rmse and that of an industrial chimney smoke cloud shown in Figure LABEL:fig:smoke_sequence. With our results, we also demonstrate how the particle filter is able to distinguish the localisation uncertainty and provide a coarse estimate of the target object’s shape and size.